OpenStudy (anonymous):

multiply and simplify in lowest terms x^2/ x^2+5* x+10/x

OpenStudy (anonymous):

\[\frac{ x^2}{ x^2+5 }*\frac{ x+10 }{ x }\]

OpenStudy (anonymous):

the x's can be cancelled

hero (hero):

Yes, correct

hero (hero):

Only in the first fraction

hero (hero):

So now, we will have: \[\frac{ x(x+10) }{ x^2+5 }\] Now you can muliply that out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x^2+10x/x^2+5?

OpenStudy (anonymous):

thanks